# CosForce: A Force-Based General Model for Simulating Pedestrian Anticipation and Reaction Mechanisms

### v1

@@ Abstract
In this study, a force-based general model, named the CosForce model, has been developed. To the best of our knowledge, this may represent the most simplified version of the social force model (streamlined rules and high efficiency). Considering the anisotropic influence of the environment on pedestrian, the model solely focuses on interactions between pedestrians and their nearest neighbors, rather than all surrounding pedestrians (as the pedestrian count increases, the computational cost transitions from power-law growth to linear growth). Specifically, cosine functions are employed to describe the interactions between nearest neighbors, allowing the model to simulate the anticipation and reaction mechanisms. By applying this model to simulations across various scenarios, a wide range of pedestrian dynamics can be simulated, including pedestrian fundamental diagrams and self-organized phenomena such as lane formation, stripe formation, and cross-channel formation, among others.
The open-source codes of the model can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link, and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
Force-based model, Pedestrian self-organization, Collective dynamics, Numerical simulation
@@ 1 Introduction
The study of pedestrian dynamics modeling has undergone several decades of development. As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic planning, currently exhibiting unprecedented vitality. Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by pedestrian interactions. The fundamental models generally consider only the real-time interaction processes between pedestrians and the environment. On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route decision in pedestrian traffic, etc. Classic approaches, such as the cellular automata model (Burstedde et al., 2001), social force model (Helbing & Molnar, 1995), reciprocal velocity obstacles model (Van den Berg et al., 2008) and the heuristic model (Moussaïd et al., 2011), remain mainstream methods.
@@
Due to their high extensibility, force-based models have been applied not only in pedestrian dynamics modeling but also in broader fields (Liu et al., 2024). New rule sets can be added by researchers to investigate specific mechanisms as required. Among these mechanisms, pedestrian anticipation and reaction have been extensively analyzed. Pedestrian reaction behaviors are generally viewed as passive responses to the environment, such as car-following behavior, whereas anticipation behaviors are considered active responses to environmental stimuli, such as a goalkeeper’s predictive dive in soccer. The subtle distinctions between anticipation and reaction mechanisms are frequently overlooked, leading to the homogenization of the modeling process. Collision avoidance rules can be introduced into force-based models in various forms, based on the relative velocity and relative distance of agents. Currently, the social force model incorporating collision avoidance rules based on the time-to-collision (TTC) estimation is commonly regarded as a model embedding anticipation mechanisms (Zanlungo et al., 2011; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
@@
A potential issue is that the addition of rule sets or new force terms may render the model cumbersome, thereby reducing its generalizability. Such models are often effective in specific scenarios, but their optimization for specific objectives may, in fact, weaken their generalization capability. In this regard, it may be more effective to develop a general model by simplifying existing models rather than by adding complexity. To this end, a model was developed to simulate anticipation and reaction mechanisms in pedestrian interaction. The rules of the model were further simplified based on the classic social force model while aiming for higher simulation accuracy.
@@ 2 Experience and Intuition in Pedestrian Interactions
In the classical social force model, pedestrians are represented as rigid particles. The volume exclusion principle, which is strictly constrained by the function $g(x)$, has been shown to potentially cause distortions in pedestrian dynamics under extreme conditions. In order to capture more detailed motion dynamics, some studies have proposed the use of ellipsoidal shapes instead of circular ones to simulate pedestrian movement (Chraibi et al., 2010). However, this modification inevitably results in increased model complexity. A more effective compromise can be achieved by considering pedestrians as elastic bodies rather than rigid ones. In narrow passages, pedestrians frequently turn sideways, and mutual compression between individuals reduces the space occupied. Therefore, modeling pedestrians as elastic bodies allows for a more accurate simulation of these behaviors.
@@
Anisotropy of pedestrian motion space
Traditionally, the space surrounding pedestrians has been assumed to be isotropic, implying that the influence exerted by pedestrians in different directions on a target pedestrian is identical. The interaction force between pedestrians has typically been modeled as an exponentially decaying function of distance. The rapid decay property ensures that force interactions are significant only between nearby pedestrians. However, this modeling approach results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations. Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009).
In real-world movement, it is generally observed that pedestrians primarily focus on the dynamics ahead of them while largely disregarding those behind them. Therefore, incorporating anisotropy in modeling pedestrian movement appears to be more consistent with the intuitive and empirically observation of pedestrians.
@@
The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems. We conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Bonnemain et al., 2023). For instance, in a typical crossing flow scenario, as depicted in Fig. 1(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate phase separation between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig. 1(b). Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig. 1(c).
Collisions are necessary
The development of models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). However, a potential issue with this model is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and movement direction are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can’t perform such calculations, so their movements rely on rough approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in high-density crowd scenarios.
3 Model
@@ 3.1 Model Rules
Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined. The HFA is constrained by the eccentricity attentional angle $\phi$ and the attentional depth $h$, as depicted in Fig. 2. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement. The attentional depth $h$ can theoretically extend to infinity. For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
Figure 2: Diagram of the process of pedestrian interaction.
In analogy to the classical Social Force Model, the pedestrian movement in the proposed model is described as a combination of a self-driving force, an inter-pedestrian interaction force, and an obstacle-pedestrian interaction force. With respect to the inter-pedestrian interaction force, due to the anisotropy of the motion space, only the force $f_{ij}$, exerted by the nearest pedestrian $j$ within the region $\boldsymbol{\Omega}$, is considered. The global constraints can be expressed as follows:
$\boldsymbol{f}={{\boldsymbol{f}}_{i}}+{{\boldsymbol{f}}_{ij}}+\sum\limits_{w}{{{\boldsymbol{f}}_{iw}}}$ (1)
@@
Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the interaction force from the nearest neighbor $j$, and $\boldsymbol{f}_{iw}$ represents the force exerted by obstacles such as walls.

### v2

# CosForce: A Force-Based General Model Considering Pedestrian Anticipation and Reaction Mechanisms

@@ Abstract
In this study, a force-based general model, named the CosForce model, has been developed. To the best of our knowledge, this may represent the most simplified version of the force-based model. Considering the anisotropic interplay of pedestrians, the model focuses on binary interactions rather than metric-based interactions. Specifically, cosine functions are employed to describe the asymmetric interactions between the focal pedestrian and their nearest neighbor, implicitly simulating anticipation and reaction mechanisms.
In addition to numerical calibration in basic scenarios, the simulation primarily focuses on two intriguing crowd phenomena: "phase separation" and the "catfish effect". Speed related metrics are introduced to quantitatively analyze self-organizing phenomena, and the "catfish effect" is examined from the perspective of momentum assessment. Many of the results require further exploration, particularly with support from empirical data. This study provides a new perspective on pedestrian modeling: the dynamics among pedestrians can be simplified as binary interactions.
agent-based models, collective dynamics, phase separation, catfish effect, numerical simulation
@@ 1 Introduction
The study of pedestrian dynamics modeling has undergone several decades of development. As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic swarms, currently exhibiting unprecedented vitality. Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by real-time pedestrian interactions. On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route choice in pedestrian traffic, etc (tactical level). Classic approaches, such as the cellular automaton model (Burstedde et al., 2001), the social force model (Helbing & Molnar, 1995), the reciprocal velocity obstacles model (Van den Berg et al., 2008), the heuristic model (Moussaïd et al., 2011), and the PLE model (Guy et al., 2010) based on Maupertuis’ Principle of Least Action, have been widely applied.
@@
Due to their high extensibility, force-based models have been applied not only in pedestrian dynamics modeling but also in broader fields (Liu et al., 2024). New rule sets can be added by researchers to investigate specific mechanisms as required. Among these mechanisms, pedestrian anticipation and reaction have been extensively analyzed. Pedestrian reaction behaviors are generally viewed as passive responses to the environment, such as car-following behavior, whereas anticipation behaviors are considered active responses to environmental stimuli, such as a goalkeeper’s predictive dive in soccer. A similar mechanism has been observed in experimental research within the field of cognitive science. Interactive experiments has revealed that visual attention is deployed differentially, depending on the nature of the behavioral goal that designates the task relevance of visual input (Renton et al., 2019; Frielink-Loing et al., 2017). The subtle distinctions between anticipation and reaction mechanisms are frequently overlooked, leading to the homogenization of the modeling process. Collision avoidance rules can be introduced into force-based models in various forms, based on the relative velocity and relative distance of agents. Currently, the model incorporating collision avoidance rules based on the time-to-collision (TTC) potential is commonly regarded as the anticipation based model (Zanlungo et al., 2011; Gerlee et al., 2017; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
@@
A potential issue is that the addition of rule sets or new force terms may render the model cumbersome, thereby reducing its generalizability. Such models are often effective in specific scenarios, but their optimization for specific objectives may, in fact, weaken their generalization capability. In this regard, it may be more effective to develop a general model by simplifying existing models rather than by adding complexity.
In consideration of applicability and stability, the general model of the operational layer should incorporate the following characteristics:
Interpretability (aligned with physical logic): dimensional consistency, scale invariance, and parameter transfer across scenarios.
Empirical foundations: grounded in empirical data and aligned with pedestrian movement intuitions.
Principles of minimalism: adherence to the simplest possible rules, the fewest parameters, and ease of analysis.
Diversity in model performance: fundamental diagrams and pedestrian self-organization, etc.
Computational efficiency: lower computational cost compared to similar models of the same type.
To this end, a general model of pedestrian dynamics is established in this paper, constructed based on binary interactions. The rules of the model, developed from first principles, were further refined in accordance with the classic social force model to enhance simulation accuracy. Additionally, this minimalist approach facilitates the analytical examination of pedestrian interactions.
@@ 2 Experience and Intuition in Pedestrian Interactions
Anisotropy of pedestrian motion space
Traditionally, the space surrounding pedestrians has been assumed to be isotropic. This method results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations. Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009). Furthermore, similar models have yet to attain optimal accuracy in crowd modeling, as they demonstrate effectiveness primarily in specific contexts, such as densely packed crowds. In such cases, vision-based or attention-driven movement patterns are ineffective, and direct physical contact leads the dense crowd to approximate the behavior of a particle system.
Empirical observations of pedestrian behavior reveal that individuals primarily focus on the dynamics in front of them, largely disregarding those behind. Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined. The HFA is constrained by the attentional eccentricity angle $\phi$ and the attentional depth $h$, as depicted in Fig. 1. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement. The attentional depth $h$ can theoretically extend to infinity. For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
Figure 1: Diagram of the process of pedestrian interaction.
@@
In the classical social force model, pedestrians are represented as rigid particles. The volume exclusion principle, which is strictly constrained by the function $g(x)$, has been shown to potentially cause distortions in pedestrian dynamics under extreme conditions. In order to capture more detailed motion dynamics, some studies have proposed the use of ellipsoidal shapes instead of circular ones to simulate pedestrian movement (Chraibi et al., 2010). On the other hand, this modification inevitably leads to an increase in model complexity. An effective compromise can be achieved by modeling pedestrians as elastic bodies rather than rigid ones. In narrow passages, pedestrians frequently turn sideways, and mutual compression between individuals reduces the space occupied. Therefore, modeling pedestrians as elastic bodies allows for a more accurate simulation of these behaviors.
@@
Collisions are necessary
The models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). However, a potential issue is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and movement direction are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can’t perform such calculations, so their movements rely on rough approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in high-density crowd scenarios.
How many pedestrians can we tracked simultaneously?
Research based on collective dynamics typically establishes models from a holistic perspective, wherein macro-level rules govern the evolution of groups, such as the BOID model (Reynolds, 1987) and network models (Bode et al., 2011; Allen et al., 2017). Based on metrics (Vicsek et al., 1995), topological (Ballerini et al., 2008), or visual connection (Rosenthal et al., 2015; Wirth et al., 2023), the state of individuals exhibits a strong correlation with the characteristics of their surrounding neighborhoods. The complexity and variability of pedestrian dynamics hinder the approximation of crowds to fish schools or flocks of birds. However, the paradigms of interactions observed can indeed be transferred across species.
When pedestrians move within a crowd, they primarily rely on vision to acquire information. The visual field of pedestrians encompasses a considerable range, potentially including dozens or even hundreds of other pedestrians and possible obstacles. However, the range within which pedestrians actually focus their attention is quite limited. The allocation of attention by pedestrians falls within the domains of cognitive sciences. In addition to the "magical number 4" theory (Cowan, 2001) concerning visual attention, mainstream perspectives suggest that the complexity and dynamism of the environment affect the number of observable entities, with higher levels of dynamism and complexity leading to a reduction in the number of objects that pedestrians can track (Alvarez & Franconeri, 2007).
In highly stochastic and dynamic crowds, pedestrians can realistically focus on only a very limited number of targets, typically the nearest neighbors within their attention field. From this perspective, models that employ binary interactions (topological connections for $n$=1) rather than metric-based interactions may be more concise and efficient.
@@
The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems. From a reductionist perspective, we conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Rand et al., 2014; Su et al., 2022; Bonnemain et al., 2023; Zablotsky et al., 2024). For instance, in a typical crossing flow scenario, as depicted in Fig. 2(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate "phase separation" between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig. 2(b). Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig. 2(c). In the model, the local cooperation rule is implicitly embedded via the asymmetric forces interaction based on the cosine constraint.
3 Methodology
3.1 Constraint of Space-speed Relationship
Figure 3: Space-speed relationship, (a) speed-local density constraint, (b) speed-headway constraint.
The relationship between speed and space has been extensively studied in the context of pedestrian dynamics, as illustrated in Fig. 2. Assuming isotropic pedestrian interactions, it has been shown that under constrained conditions, velocity is inversely proportional to local density (i.e., the speed–local density relationship), as expressed in Eq. 1, given a constant parameter of $k$. Based on this empirical relationship, within the framework of force-based models, the variation in pedestrian force with respect to spatial metrics can be described by a power function ($\partial t=\tau$), as explained in Eq. 2.
$v\propto\rho^{-1}\propto d^{2}$ (1)
$\partial v&=k\partial d^{2}\;(\text{omit}\;{\partial t}),\\ \displaystyle\partial f&=m\frac{\partial v}{\partial t}=mk\frac{\partial d^{2}}{\partial t}.\end{aligned}\right.$ (2)
Furthermore, in one-dimensional motion, a linear empirical relationship between the headway and speed is typically employed, often associated with the principle of constant time headway. Considering the anisotropy of the pedestrian movement space, this linear relationship, based on the nearest neighbor distance, can similarly be extended to two-dimensional motion, as illustrated in Eq. 3. In this context, a linear function can be employed to describe the variation of pedestrian force with respect to nearest neighbor relative distance ($\partial t=\tau$), as shown in Eq. 4.
$d=vt_{h}+d_{0}\quad\Rightarrow\quad v=\frac{d-d_{0}}{t_{h}}$ (3)
$\frac{\partial d}{\partial v}&=t_{h}\;(\text{omit}\;{\partial t}),\\ \partial f&=\frac{m\partial v}{\partial t}=\frac{m\partial d}{t_{h}\,\partial t}.\end{aligned}\right.$ (4)
Within the framework of force-based models, based on these property, it is natural to employ power functions or linear functions of distance to describe such mechanisms. The empirical formulas underlying linear functions demonstrates dimensional consistency and scale invariance, contributes to its stability.
Based on the hypothesis of binary interactions, each pedestrian in a non-collision condition both subject to only two forces: the self-driving force and the nearest-neighbor repulsive force. Consequently, the equilibrium properties of pedestrians in a one-dimensional scenario can be derived as follows:
$v\to\max\left(\min\left(\frac{{d_{ij}-r_{ij}}}{{t_{h}}},v_{\max}\right),0\right)\Rightarrow f_{ij}\to f_{i}.$ (5)
The scalar form of the self-driving force $f_{i}$ is defined by Eq. 6. Therefore, the equation for the nearest-neighbor repulsive force at equilibrium can be expressed as shown in Eq. 7.
${f}_{i}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}-{v}_{i}}\right)$ (6)
${f}_{ij}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}}-\max\left({\min\left({\frac{{{{d}_{ij}}-{r_{ij}}}}{{{t_{h}}}},{{v}_{\max}}}\right),0}\right)\right)$ (7)
3.2 Anticipation and reaction mechanisms within force-based model framework
[…]
In the computation based on two-dimensional TTC, the angle $\theta$ between the relative velocity and relative distance plays a critical role in pedestrians’ collision avoidance processes ($\cos\theta$ represents the projection scale between the relative velocity and relative distance). An additional scaling factor ($1+\alpha\cos\theta$) is introduced into the nearest-neighbor repulsive force to describe the effect of collision avoidance. Here, $\alpha$ is a dimensionless coefficient that regulates the collision avoidance scale of pedestrians, satisfying $(1+\alpha\cos\theta)\in[1-\alpha,1+\alpha]$. $\alpha\to 0$ indicates that pedestrians are insensitive to collisions, which typically corresponds to unidirectional flow scenarios. $\alpha\to 1$ signifies that pedestrians are highly sensitive to collisions, corresponding to crowd or multiphase flow scenarios.
Symmetric and Aymmetric forces
Based on the HFA described above, pedestrian interactions involve both reciprocal and non-reciprocal interactions, as illustrated in Fig. 4. Symmetric and asymmetric forces naturally be employed to represent these mechanisms. Overall, these mechanisms are associated with the anticipation and reaction behaviors observed in pedestrian dynamics. In this context, the anticipation mechanism is referred to as a implicit cooperative game process among conflicting pedestrians (inducing separation). Conversely, the reaction mechanism is described as the convergence process among co-directional pedestrians (inducing aggregation).
3.2.1 Individual-level analysis
The nearest-neighbor interaction force $\boldsymbol{f}_{ij}$ is adjusted nonlinearly using a cosine function of the angle $\theta$. When $\theta\to 0$, it indicates a potential collision trend, and $\boldsymbol{f}_{ij}$ increases. When $\theta\to\pi$, it indicates a separation trend between pedestrians, and $\boldsymbol{f}_{ij}$ decreases. With the introduction of the cosine function, the phenomena of separation and aggregation have been implicitly simulated.
Separation of conflicting flows
Fig. 4(c) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces. Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,x}\right\|<\left\|\boldsymbol{f}_{ij,y}\right\|\right)$. This trend is gradually amplified with the update of their relative positions, ultimately leading to agent $i$ yielding to agent $j$ (cooperative game). Strictly speaking, if the agents’ radii are ignored, non-cooperative games occur only when $\theta=0$.
Aggregation of co-directional flows
Fig. 4(d) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$. If $\left\|\boldsymbol{v}_{j}\right\|>\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ increases, and $\boldsymbol{v}_{i}$ decreases. If $\left\|\boldsymbol{v}_{j}\right\|<\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ decreases, and $\boldsymbol{v}_{i}$ increases (considering the self-driving force). This leads to the aggregation of pedestrians in the same direction. Meanwhile, it satisfies the speed-density relationship under the constraints of the fundamental diagram.
3.2.2 Analytical Analysis of the Nearest Neighbor Repulsive Force
Based on the aforementioned deduce, we derived the equation of the nearest-neighbor repulsive force as a function of distance $d$ and angle $\theta$, as shown in Eq. 8.
${f}_{ij}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}}-\max\left({\min\left({\frac{{{{d}_{ij}}-{r_{ij}}}}{{{t_{h}}}},{{v}_{\max}}}\right),0}\right)\right){\left({1+\alpha\cos\theta}\right)}$ (8)
For the convenience of differential computation, the tanh function can be employed to describe its continuous form, which aligns with the formulation used to represent the desired velocity in traffic flow. The continuous form is provided in Eq. 9.
${f}_{ij}=\frac{{m_{i}}}{\tau}\left({v}_{\max}\left(1-\tanh\left(\frac{{d}_{ij}-r_{ij}}{t_{h}}\right)\right)\right)\left(1+\alpha\cos\theta\right)$ (9)
Fig. 5 presents the influence of the nearest-neighbor relative distance $d$ and angle $\theta$ on the force term. Empirical results can be used for surface fitting to investigate a more specific relationship, which may not follow a cosine function.
Figure 5: Relationship between the repulsive force and $d$ and $\theta$, ($\alpha=0.5$), above: piecewise function, below: continuous function.
@@ 3.3 Model Rules
Based on the framework of a force-based model, the model is characterized by a self-driving force, a repulsive force derived from the space-speed constraint, and the contact forces that arise during collisions. Based on the assumption of binary interaction, only the force $\boldsymbol{f}_{ij}$, exerted by the nearest entity (pedestrian or wall) $j$ ($j\in\boldsymbol{\Omega}:=\left\|{{\boldsymbol{d}_{ij}}}\right\|<h\wedge\angle\left({{\boldsymbol{v}_{i}},{\boldsymbol{n}_{ij}}}\right)<\phi$), is considered. The global constraints can be expressed as follows:
$\boldsymbol{f}=\boldsymbol{f}_{i}+\boldsymbol{f}_{ij}+\sum_{\|\boldsymbol{d}_{ij}\|<r_{ij}}\boldsymbol{f}_{c}=m\boldsymbol{a}$ (10)
@@
Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the repulsive force from the nearest neighbor $j$.